OpenStudy (anonymous):

A solid cone of radius 6 cm and height 21 cm is melted and made into 2 spheres of different sizes.The radius of the first sphere is 4 cm , find the radius of the second sphere.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: the volume of the 2 spheres has to be equal to the volume of the starting cone, so we can write this equation: \[\Large \frac{{4\pi }}{3}{R^3} + \frac{{4\pi }}{3}{x^3} = \frac{{\pi {r^2}h}}{3}\] where x is the requested radius, R=4 cm, h=21 cm, and r= 6 cm

OpenStudy (anonymous):

???

OpenStudy (michele_laino):

please you have to solve that equation for x

OpenStudy (anonymous):

i didnt how you got yhe equation *the

OpenStudy (michele_laino):

the sum of the volumes of the 2 spheres, has to be equal to the volume of the starting cone

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\pi (R^3 + r^3) = \frac{ 1 }{ 3 }\pi(r^2h)\]

OpenStudy (michele_laino):

not exactly, better is: \[\Large \frac{{4\pi }}{3}\left( {{R^3} + {x^3}} \right) = \frac{{\pi {r^2}h}}{3}\] where x is the requested radius, and r is the radius of the starting cone, namely r= 6 cm

OpenStudy (anonymous):

3 and pi get cancelled?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

r^h = 4(R^3 + x^3)?

OpenStudy (michele_laino):

more precisely: \[\Large 4\left( {{R^3} + {x^3}} \right) = {r^2}h\]

OpenStudy (anonymous):

what next?

OpenStudy (michele_laino):

we have to divide both sides by 4

OpenStudy (anonymous):

(R^3 +x^3) = r^2h/4

OpenStudy (michele_laino):

ok! now we have subtract R^3 at both sides

OpenStudy (anonymous):

x^3 = r^2h/4 - R^3

OpenStudy (michele_laino):

finally we have to take the 3-rd root of both sides

OpenStudy (anonymous):

???

OpenStudy (michele_laino):

you should get this: \[\Large x = \sqrt[3]{{\frac{{{r^2}h}}{4} - {R^3}}}\]

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

\[\Large x = \sqrt[3]{{\frac{{{r^2}h}}{4} - {R^3}}} = \sqrt[3]{{\frac{{{6^2} \cdot 21}}{4} - {4^3}}} = ...?\]

OpenStudy (anonymous):

5?

OpenStudy (michele_laino):

that's right! x= 5 cm

OpenStudy (anonymous):

thanks.

OpenStudy (michele_laino):

:)